OpenStudy (anonymous):

evaluate the following expression using the values given: -a2 - 3b3 + c2 + 2b3 - c2 if a = 3, b = 2, and c = -3

OpenStudy (anonymous):

\[-a ^{2}-3b ^{3}+c ^{2}-2b ^{3}-c ^{2}\]

OpenStudy (anonymous):

is that the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[-(3)^2 -3(2)^3 + (-3)^2 - 2(2)^3 - (-3)^2\] -9-24+9-16-9 -49

OpenStudy (anonymous):

thanx😊

OpenStudy (aum):

I would simplify the expression first and then substitute: \(\large -a^2 - 3b^3 + c^2 + 2b^3 - c^2 = -a^2 - b^3\) Put a = 3 and b = 2 \(\large -3^2 - 2^3 = -9 - 8 = -17\)